OpenStudy (anonymous):

I need help writing a sine function Amp:4 period 3(pi) horizontal shift 1 unit right vertical shift 6units up

OpenStudy (anonymous):

The basic sine function form is \(f(x)=a\sin(k(x-b))+c\). Amplitude being 4 implies a = 4. Period being \(3\pi\) implies \[3\pi=\frac{2\pi}{k}\text{ then solve for k.}\] The horizontal shift 1 unit to the right implies b = +1 (positive because to the right.) Vertical shift 6 units up implies c = +6 (positive because upward.)

OpenStudy (anonymous):

Ok great thank you

OpenStudy (anonymous):

your welcome. :)

OpenStudy (anonymous):

Would K be 2/3

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

So it would look like f(x) = 4sin (2/3x+1))+6

OpenStudy (anonymous):

Nope. It should be \(f(x)=4\sin\left(\frac{2}{3}(x-(+1))\right)+(+6)\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

:))